# The ALMA Survey of Gas Evolution of PROtoplanetary Disks (AGE-PRO): I. Program Overview and Summary of First Results

### v1

@@ 6.3.3 Theoretical interpretation of the AGE-PRO results
AGE-PRO VI (Kurtovic et al., 2025, in press) compares the mm continuum fluxes, dust disk sizes, and spectral indexes observed in the AGE-PRO sample with dust evolution simulations. They find that the AGE-PRO results are consistent with models with weak or strong pressure traps. The evolution of the spectral index in the AGE-PRO sample is also suggestive of an unresolved population of dust traps.
AGE-PRO VII (Tabone et al., 2025, in press) develops a population synthesis approach to test two scenarios for disk evolution: MHD disk-wind driven and turbulence-driven with internal photoevaporation winds. A systematic method was used to fit the distribution of four disk properties of Class II disks: disk fractions, mass accretion rates, gas disk masses and sizes. For Class I/FS sources, only the gas masses are used for constraining models due to their large uncertainties in accretion rates and gas disk sizes. The results show that the MHD wind-driven accretion can reproduce the bulk disk properties from Ophiuchus to Upper Sco, assuming initial compact disks with a magnetic field that declines with time. In contrast, the best-fitting models of turbulent-driven disks with internal photoevaporation wind overpredict the median disk masses in the Lupus and Upper Sco regions by one order of magnitude.
@@ 6.3.4 Interesting individual disks
AGE-PRO XII (Miley et al., 2025, in press) reports a flare detected in the continuum emission of the 2MASS J16202 863-2442087 (Upper Sco 7 in AGE-PRO) system, which decreases by a factor of 8 in less than an hour, and by a factor of 13 within 8 days. In contrast, the continuum emission of other four Upper Sco sources observed within the same execution blocks were consistent within 5% among different execution blocks, suggesting the variation of Upper Sco 7 was real. Miley et al. (2025, in press) discusses various possibilities of the flare and suggests that synchrotron emission was a probable cause of the continuum flare.
@@ 7.1.1 Similarity of initial and environmental conditions
Our knowledge of older star-forming regions ($>$3 Myr) is also limited. Upper Sco has been the only reference point of old star-forming regions. Although the $\sigma$-Orionis region is also considered relatively old (3-5 Myr), a large fraction of its disk population is under a strong external FUV field ($>$1000 G0, Ansdell et al., 2017), which causes strong external photoevaporation and significantly affects disk evolution (Maucó et al., 2023). In the Upper Sco region the average external FUV field is more modest ($\sim$10 G0), but is still significantly higher than that of Taurus and Lupus (a few G0). The modest FUV strength can still affect the size of disks and the disk mass to a lesser degrees. See detailed discussions on the effect of external UV on Upper Sco disks in AGE-PRO VIII (Anania et al., 2025, in press). In the future, studies of more star-forming regions are needed to understand the similarities and differences among different regions.
@@ 7.1.2 Sample selection biases
In Section 2, we discussed the selection criteria of the AGE-PRO sample. The sample was selected mainly based on spectral type, disk class type, and mm-luminosity. For the Lupus and Upper Sco regions, K-S tests of mm-luminosity and accretion rate between the AGE-PRO sample and the full population of Class II disks around M3-K6 stars showed that these are statistically indistinguishable (Deng et al., 2025, in press; Agurto-Gangas et al., 2025, in press). The representativeness of the Ophiuchus sample is more challenging to quantify, as the spectral types and mass accretion rates of the full sample of embedded disks are unknown due to high extinctions. One particular surprise of the Ophiuchus sample is that the majority of these disks are highly-inclined ($\gtrsim$ 60 deg, Ruiz-Rodriguez et al. 2025, in press; Vioque et al. 2025, in press), at a much higher proportion than the chance of random selection of inclination. Since most of the AGE-PRO Ophiuchus disks were not spatially resolved in previous surveys, only the new AGE-PRO observations revealed that these disks are highly-inclined. One possible explanation is that we selected Class I/FS sources in the Ophiuchus region from previous classifications that were based on their shallow SED slope at near-IR wavelengths. But highly-inclined disks are known to produce shallow SED near-IR slope. Please see Ruiz-Rodriguez et al. (2025, in press) for in-depth discussion on other indicators of the youth of the Ophiuchus sample.
An inevitable bias for surveys to study properties in old disks is the survivor bias, because observations can only be done for the disk population that have survived up to the given age. Almost all mm disk surveys selected samples from sources with IR excesses and/or accretion signatures (e.g., Ansdell et al., 2016; Pascucci et al., 2016; Barenfeld et al., 2016; Cieza et al., 2019). The disk fraction with these features is known to decrease with age, from $\sim$80-90in $<$1 Myr-old regions to 5-10% in 5-10 Myr-old regions, with an exponentially decrease timescale of 2-3 Myr (e.g., Haisch et al., 2001; Fedele et al., 2010; Ribas et al., 2014). As a result, the median disk properties are the median properties of the survivors. Therefore, the evolution of the median properties cannot be directly interpreted as the general evolutionary pathway of an individual disk. AGE-PRO VII (Tabone et al., 2025, in press) combines disk fractions, mass accretion rates, and the AGE-PRO results of disk masses and sizes, in a population analysis approach, using these constraints together to test both turbulent viscous and MHD evolution models.
@@ 7.2 Decoupled evolution in gas and dust disk masses
The AGE-PRO program is the first systematic study to observationally trace gas disk mass evolution in protoplanetary disks. The new constraints on gas disk masses provide evidence that the evolution of dust and gas disk masses probably does not operate on the same timescale. The different evolution timescale between gas and dust masses leads to a surprising swing in the gas-to-dust mass ratio, with a median ratio first decreasing and then increasing. These differences confirm that dust masses cannot be used to obtain reliable gas disk masses. AGE-PRO VII (Tabone et al., 2025, in press) compares two gas evolution mechanisms: viscosity and MHD-disk driven, and discusses how AGE-PRO results constrain the timescales of these mechanisms.
@@ 7.3 The typical gas and dust disk sizes
The sizes of gas and dust disks provide important insight into the disk evolution. As shown in Table 4, the median dust disk sizes are remarkably similar among the three samples, from 30-42 au. The possible lack of evolution in the dust disk sizes suggests that some level of pressure traps must be present to keep some mm-sized particles in the outer region of the disks. On the other hand, the decreases in dust masses over time suggest that pressure traps in the outer disk regions cannot completely block the inward pebble drift. Please see Kurtovic et al. (2025, in press) for an in-depth discussion on the level of pressure traps needed to explain the dust disk size and mm continuum fluxes.
@@
The apparent lack of gas disk size evolution might seem to disprove the theory of viscous evolution, as viscously evolving disks are expected to expand over time (Trapman et al., 2020b). However, both internal and external photoevaporation can cause gas disk sizes to shrink over time (Tabone et al., 2025, in press; Anania et al., 2025, in press). Therefore, gas disk size alone is not decisive evidence against viscous evolution. Nonetheless, it remains a critical input for population synthesis models to constrain disk evolution mechanisms (Tabone et al., 2025, in press).
@@ 7.4 CO gas abundance evolution
In the AGE-PRO study, we see a wide range of CO gas abundances ($x_{\rm CO}$) between $10^{-6}$ and $10^{-4}$ in the Lupus and Upper Sco disks (Trapman et al., in press, 2025a). However, the majority of our disks have $x_{\rm CO}$ around $10^{-5}$. These results support significant CO depletion in 1-6 Myr disks, but few extreme cases like $x_{\rm CO}\sim 10^{-6}$ observed in the TW Hya disk. Notably, we do not observe significant variation in $x_{\rm CO}$ between the Lupus and Upper Sco sources, suggesting that CO depletion generally ceases to progress efficiently after the first 1-3 Myr.
@@ 8 Conclusion and Future directions
The AGE-PRO program measures the gas and dust disk masses and sizes for 30 disks around M3-K6 stars in three star-forming regions, including the Ophiuchus sample ($<$1 Myr), the Lupus sample (1-3 Myr), and the Upper Sco sample (2-6 Myr). The results provide the first systematic sample to study gas evolution in protoplanetary disks. Our main findings are:
@@
    External photoevaporation can significantly reduce gas disk sizes, even under moderate far-ultraviolet (G0$\sim$10) radiation fields. While it also decreases gas disk masses, the effect is less pronounced than for disk sizes.
@@
This paper makes use of the following ALMA data: ADS/JAO.ALMA#2021.1.00128.L. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. We thank the anonymous referee whose comments helped us improve the manuscript. We thank Alena Rottensteiner and Phil Armitage for the useful discussions that contributed to this work. K.Z. and L.T. acknowledge the support of the NSF AAG grant #2205617. G.R. acknowledges funding from the Fondazione Cariplo, grant no. 2022-1217, and the European Research Council (ERC) under the European Union’s Horizon Europe Research & Innovation Programme under grant agreement no. 101039651 (DiscEvol). Views and opinions expressed are however those of the author(s) only, and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. P.P. and A.S. acknowledge the support from the UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee from ERC (under grant agreement No 101076489). A.S. also acknowledges support from FONDECYT de Postdoctorado 2022 ##\##3220495. B.T. acknowledges support from the Programme National “Physique et Chimie du Milieu Interstellaire” (PCMI) of CNRS/INSU with INC/INP and co-funded by CNES. I.P. and D.D. acknowledge support from Collaborative NSF Astronomy & Astrophysics Research grant (ID: 2205870). C.A.G. and L.P. acknowledge support from FONDECYT de Postdoctorado 2021 #3210520. L.P. also acknowledges support from ANID BASAL project FB210003. L.A.C and C.G.R. acknowledge support from the Millennium Nucleus on Young Exoplanets and their Moons (YEMS), ANID - Center Code NCN2021_080 and L.A.C. also acknowledges support from the FONDECYT grant #1241056. N.T.K. acknowledges support provided by the Alexander von Humboldt Foundation in the framework of the Sofja Kovalevskaja Award endowed by the Federal Ministry of Education and Research. K.S. acknowledges support from the European Research Council under the Horizon 2020 Framework Program via the ERC Advanced Grant Origins 83 24 28. E.T. acknowledges support from the National Science Foundation Graduate Research Fellowship program.
@@ Appendix A SEDs
Figure 9: Spectral Energy Distributions (SED) of AGE-PRO sources. Blue curves show the PHOENIX stellar photosphere models using the stellar parameters and extinction values listed in Table 2.1. The red curves are archival Spitzer IRS spectra. These SEDs and detailed references are shown in Ruiz-Rodriguez et al. (2025, in press); Deng et al. (2025, in press); Agurto-Gangas et al. (2025, in press).
@@ Appendix B Ages of the Lupus and Upper Sco sources in AGE-PRO
Figure 10: Isochronal ages of the Lupus and Upper Sco samples of AGE-PRO. The circles are individual ages and the squares represent the median ages of each group. The ages were derived based on stellar properties listed in Table 2.1 and PMS evolutionary tracks of Feiden (2016); Baraffe et al. (2015). The individual ages and their uncertainties can be found in Deng et al. (2025, in press); Agurto-Gangas et al. (2025, in press).
Figure 11: Comparison between the ages derived by the AGE-PRO program (Deng et al., 2025, in press; Agurto-Gangas et al., 2025, in press) and isochronal ages from CMD fitting of astrometric clusters by Ratzenböck et al. (2023a, b). Although the ages of Ratzenböck et al. (2023a, b) are older than our results, the median ages of Upper Sco sources are consistently older than that of Lupus sources in both age estimations.

### v2

@@ 6.3.3 Theoretical interpretation of the AGE-PRO results
AGE-PRO VI (Kurtovic et al., 2025) compares the mm continuum fluxes, dust disk sizes, and spectral indexes observed in the AGE-PRO sample with dust evolution simulations. They find that the AGE-PRO results are consistent with models with weak or strong pressure traps. The evolution of the spectral index in the AGE-PRO sample is also suggestive of an unresolved population of dust traps.
AGE-PRO VII (Tabone et al., 2025) develops a population synthesis approach to test two scenarios for disk evolution: MHD disk-wind driven and turbulence-driven with internal photoevaporation winds. A systematic method was used to fit the distribution of four disk properties of Class II disks: disk fractions, mass accretion rates, gas disk masses and sizes. For Class I/FS sources, only the gas masses are used for constraining models due to their large uncertainties in accretion rates and gas disk sizes. The results show that the MHD wind-driven accretion can reproduce the bulk disk properties from Ophiuchus to Upper Sco, assuming initial compact disks with a magnetic field that declines with time. In contrast, the best-fitting models of turbulent-driven disks with internal photoevaporation wind overpredict the median disk masses in the Lupus and Upper Sco regions by one order of magnitude.
@@ 6.3.4 Interesting individual disks
AGE-PRO XII (Miley et al., 2025) reports a flare detected in the continuum emission of the 2MASS J16202 863-2442087 (Upper Sco 7 in AGE-PRO) system, which decreases by a factor of 8 in less than an hour, and by a factor of 13 within 8 days. In contrast, the continuum emission of other four Upper Sco sources observed within the same execution blocks were consistent within 5% among different execution blocks, suggesting the variation of Upper Sco 7 was real. Miley et al. (2025) discusses various possibilities of the flare and suggests that synchrotron emission was a probable cause of the continuum flare.
@@ 7.1.1 Similarity of initial and environmental conditions
Our knowledge of older star-forming regions (>>3 Myr) is also limited. Upper Sco has been the only reference point of old star-forming regions. Although the $\sigma$-Orionis region is also considered relatively old (3-5 Myr), a large fraction of its disk population is under a strong external FUV field (>>1000 G0, Ansdell et al., 2017), which causes strong external photoevaporation and significantly affects disk evolution (Maucó et al., 2023). In the Upper Sco region the average external FUV field is more modest ($\sim$10 G0), but is still significantly higher than that of Taurus and Lupus (a few G0). The modest FUV strength can still affect the size of disks and the disk mass to a lesser degrees. See detailed discussions on the effect of external UV on Upper Sco disks in AGE-PRO VIII (Anania et al., 2025). In the future, studies of more star-forming regions are needed to understand the similarities and differences among different regions.
@@ 7.1.2 Sample selection biases
In Section 2, we discussed the selection criteria of the AGE-PRO sample. The sample was selected mainly based on spectral type, disk class type, and mm-luminosity. For the Lupus and Upper Sco regions, K-S tests of mm-luminosity and accretion rate between the AGE-PRO sample and the full population of Class II disks around M3-K6 stars showed that these are statistically indistinguishable (Deng et al., 2025; Agurto-Gangas et al., 2025). The representativeness of the Ophiuchus sample is more challenging to quantify, as the spectral types and mass accretion rates of the full sample of embedded disks are unknown due to high extinctions. One particular surprise of the Ophiuchus sample is that the majority of these disks are highly-inclined ($\gtrsim$ 60 deg, Ruiz-Rodriguez et al. 2025; Vioque et al. 2025), at a much higher proportion than the chance of random selection of inclination. Since most of the AGE-PRO Ophiuchus disks were not spatially resolved in previous surveys, only the new AGE-PRO observations revealed that these disks are highly-inclined. One possible explanation is that we selected Class I/FS sources in the Ophiuchus region from previous classifications that were based on their shallow SED slope at near-IR wavelengths. But highly-inclined disks are known to produce shallow SED near-IR slope. Please see Ruiz-Rodriguez et al. (2025) for in-depth discussion on other indicators of the youth of the Ophiuchus sample.
An inevitable bias for surveys to study properties in old disks is the survivor bias, because observations can only be done for the disk population that have survived up to the given age. Almost all mm disk surveys selected samples from sources with IR excesses and/or accretion signatures (e.g., Ansdell et al., 2016; Pascucci et al., 2016; Barenfeld et al., 2016; Cieza et al., 2019). The disk fraction with these features is known to decrease with age, from $\sim$80-90% in <<1 Myr-old regions to 5-10% in 5-10 Myr-old regions, with an exponentially decrease timescale of 2-3 Myr (e.g., Haisch et al., 2001; Fedele et al., 2010; Ribas et al., 2014). As a result, the median disk properties are the median properties of the survivors. Therefore, the evolution of the median properties cannot be directly interpreted as the general evolutionary pathway of an individual disk. AGE-PRO VII (Tabone et al., 2025) combines disk fractions, mass accretion rates, and the AGE-PRO results of disk masses and sizes, in a population analysis approach, using these constraints together to test both turbulent viscous and MHD evolution models.
@@ 7.2 Decoupled evolution in gas and dust disk masses
The AGE-PRO program is the first systematic study to observationally trace gas disk mass evolution in protoplanetary disks. The new constraints on gas disk masses provide evidence that the evolution of dust and gas disk masses probably does not operate on the same timescale. The different evolution timescale between gas and dust masses leads to a surprising swing in the gas-to-dust mass ratio, with a median ratio first decreasing and then increasing. These differences confirm that dust masses cannot be used to obtain reliable gas disk masses. AGE-PRO VII (Tabone et al., 2025) compares two gas evolution mechanisms: viscosity and MHD-disk driven, and discusses how AGE-PRO results constrain the timescales of these mechanisms.
@@ 7.3 The typical gas and dust disk sizes
The sizes of gas and dust disks provide important insight into the disk evolution. As shown in Table 4, the median dust disk sizes are remarkably similar among the three samples, from 30-42 au. The possible lack of evolution in the dust disk sizes suggests that some level of pressure traps must be present to keep some mm-sized particles in the outer region of the disks. On the other hand, the decreases in dust masses over time suggest that pressure traps in the outer disk regions cannot completely block the inward pebble drift. Please see Kurtovic et al. (2025) for an in-depth discussion on the level of pressure traps needed to explain the dust disk size and mm continuum fluxes.
@@
The apparent lack of gas disk size evolution might seem to disprove the theory of viscous evolution, as viscously evolving disks are expected to expand over time (Trapman et al., 2020b). However, both internal and external photoevaporation can cause gas disk sizes to shrink over time (Tabone et al., 2025; Anania et al., 2025). Therefore, gas disk size alone is not decisive evidence against viscous evolution. Nonetheless, it remains a critical input for population synthesis models to constrain disk evolution mechanisms (Tabone et al., 2025).
@@ 7.4 CO gas abundance evolution
In the AGE-PRO study, we see a wide range of CO gas abundances ($x_{\rm CO}$) between $10^{-6}$ and $10^{-4}$ in the Lupus and Upper Sco disks (Trapman et al., 2025a). However, the majority of our disks have $x_{\rm CO}$ around $10^{-5}$. These results support significant CO depletion in 1-6 Myr disks, but few extreme cases like $x_{\rm CO}\sim 10^{-6}$ observed in the TW Hya disk. Notably, we do not observe significant variation in $x_{\rm CO}$ between the Lupus and Upper Sco sources, suggesting that CO depletion generally ceases to progress efficiently after the first 1-3 Myr.
@@ 8 Conclusion and Future directions
The AGE-PRO program measures the gas and dust disk masses and sizes for 30 disks around M3-K6 stars in three star-forming regions, including the Ophiuchus sample (<<1 Myr), the Lupus sample (1-3 Myr), and the Upper Sco sample (2-6 Myr). The results provide the first systematic sample to study gas evolution in protoplanetary disks. Our main findings are:
@@
[…]
@@
This paper makes use of the following ALMA data: ADS/JAO.ALMA#2021.1.00128.L,
ADS/JAO.ALMA#2019.A.00034.S,
ADS/JAO.ALMA#2017.1.00569.S,
ADS/JAO.ALMA#2019.1.01135.S,
ADS/JAO.ALMA#2015.1.01199.S.
ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. We thank the anonymous referee whose comments helped us improve the manuscript. We thank Alena Rottensteiner and Phil Armitage for the useful discussions that contributed to this work. K.Z. and L.T. acknowledge the support of the NSF AAG grant #2205617. G.R. acknowledges funding from the Fondazione Cariplo, grant no. 2022-1217, and the European Research Council (ERC) under the European Union’s Horizon Europe Research & Innovation Programme under grant agreement no. 101039651 (DiscEvol). Views and opinions expressed are however those of the author(s) only, and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. P.P. and A.S. acknowledge the support from the UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee from ERC (under grant agreement No 101076489). A.S. also acknowledges support from FONDECYT de Postdoctorado 2022 #\#3220495. B.T. acknowledges support from the Programme National “Physique et Chimie du Milieu Interstellaire” (PCMI) of CNRS/INSU with INC/INP and co-funded by CNES. I.P. and D.D. acknowledge support from Collaborative NSF Astronomy & Astrophysics Research grant (ID: 2205870). C.A.G. and L.P. acknowledge support from FONDECYT de Postdoctorado 2021 #3210520. L.P. also acknowledges support from ANID BASAL project FB210003. L.A.C and C.G.R. acknowledge support from the Millennium Nucleus on Young Exoplanets and their Moons (YEMS), ANID - Center Code NCN2024_001 and L.A.C. also acknowledges support from the FONDECYT grant #1241056. N.T.K. acknowledges support provided by the Alexander von Humboldt Foundation in the framework of the Sofja Kovalevskaja Award endowed by the Federal Ministry of Education and Research. K.S. acknowledges support from the European Research Council under the Horizon 2020 Framework Program via the ERC Advanced Grant Origins 83 24 28. E.T. acknowledges support from the National Science Foundation Graduate Research Fellowship program.
@@ Appendix A SEDs
Figure 9: Spectral Energy Distributions (SED) of AGE-PRO sources. Blue curves show the PHOENIX stellar photosphere models using the stellar parameters and extinction values listed in Table 2.1. The red curves are archival Spitzer IRS spectra. These SEDs and detailed references are shown in Ruiz-Rodriguez et al. (2025); Deng et al. (2025); Agurto-Gangas et al. (2025).
@@ Appendix B Ages of the Lupus and Upper Sco sources in AGE-PRO
Figure 10: Isochronal ages of the Lupus and Upper Sco samples of AGE-PRO. The circles are individual ages and the squares represent the median ages of each group. The ages were derived based on stellar properties listed in Table 2.1 and PMS evolutionary tracks of Feiden (2016); Baraffe et al. (2015). The individual ages and their uncertainties can be found in Deng et al. (2025); Agurto-Gangas et al. (2025).
Figure 11: Comparison between the ages derived by the AGE-PRO program (Deng et al., 2025; Agurto-Gangas et al., 2025) and isochronal ages from CMD fitting of astrometric clusters by Ratzenböck et al. (2023a, b). Although the ages of Ratzenböck et al. (2023a, b) are older than our results, the median ages of Upper Sco sources are consistently older than that of Lupus sources in both age estimations.